OpenStudy (anonymous):

What is the volume of a square prism whose base edges are 2 centimeters and height is 8 centimeters?

OpenStudy (ash2326):

let's find the area of the base which is 2*2= 4 cm^2 volume of square prism= area of base* height = 4*8=32 cm^3